# UniBERT: Adversarial Training for Language-Universal Representations

### v2

@@ 1 Introduction
Natural language processing (NLP) has significantly advanced in recent years, primarily due to breakthroughs in deep learning and Transformer-based architectures [1]. Models such as Bidirectional Encoder Representations from Transformers (BERT) [2] and its successors [3, 4] have redefined the state of the art in tasks such as machine translation, sentiment analysis, summarization, and question answering by learning contextualized representations that capture both syntactic and semantic nuances. This evolution has not only elevated task-specific performance, but also opened new avenues for exploring the complexities of human language.
@@
At the same time, there has been a surge of interest in developing multilingual systems capable of processing and understanding multiple languages simultaneously [5, 6, 7]. Multilingual training leverages shared linguistic features across diverse languages, a benefit that is especially pronounced in low-resource settings. In addition, adversarial training [8] has emerged as a powerful mechanism to encourage the learning of language-invariant representations, thus reducing language-specific biases and improving cross-lingual generalization in a variety of tasks [9, 10].
@@
However, developing robust multilingual models poses several challenges, including managing heterogeneous data distributions, accommodating language-specific nuances, and needing scalability without compromising performance. Adversarial training addresses these difficulties by introducing a competitive dynamic that forces the model to focus on features common to all languages. This approach not only boosts overall performance but also diminishes the model’s reliance on language-specific signals to enable better generalization across many linguistic domains [11].
@@
In this work, we introduce UniBERT, a compact and efficient BERT-based model specifically designed for multilingual processing, available in three open-source versions: (1) UniBERT-Small [^1] https://huggingface.co/avramandrei/unibert-small, (2) UniBERT-XSmall [^2] https://huggingface.co/avramandrei/unibert-xsmall, and (3) UniBERT-XXSmall [^3] https://huggingface.co/avramandrei/unibert-xxsmall. UniBERT was pre-trained on a curated Wikipedia corpus spanning 107 languages and combines three key training components: masked language modeling [2], adversarial training, and knowledge distillation [12]. This integrated methodology enables the model to achieve robust performance while significantly reducing the parameter count relative to larger models, making it an attractive solution for resource-constrained applications.
@@
We perform extensive evaluations of UniBERT on four NLP tasks: (1) named entity recognition, (2) natural language inference, (3) question answering, and (4) semantic textual similarity. The results demonstrate that our models, fine-tuned jointly on all languages, obtain better results than when they are fine-tuned on each language individually. Concretely, on average, UniBERT-Small improves the performance from 47.10% to 48.87%, UniBERT-XSmall from 40.20% to 44.35%, and UniBERT-XXSmall from 34.51% to 37.64% when trained jointly in all languages, yielding an average relative improvement of 7.72%. In contrast, larger language models show an improvement of 1.17%, underscoring the significant benefits of the joint multilingual training approach enhanced by adversarial objectives.
@@
We also performed a comprehensive statistical comparative analysis to confirm the strengths of our approach. The student’s t-test yielded a p-value of 0.0181, demonstrating that gains in UniBERT performance are significant at a significance threshold of $\alpha=0.05$. It also shows that the combination of joint multilingual training and adversarial training effectively compensates for the reduced capacity of compact models.
@@
  • We show that joint multilingual training, combined with an adversarial objective, significantly enhances cross-lingual performance, particularly in resource-constrained settings.
@@ 2.1 Multilingual Language Models
The development of multilingual sentence embeddings has evolved from initial methods such as LASER [13], which leveraged a shared long short-term memory [14] encoder-decoder framework to align sentence representations across languages using parallel data. Subsequent models, such as SBERT-distill [15], utilized knowledge distillation to transfer monolingual sentence embedding capabilities into a multilingual space, allowing robust performance with fewer parallel sentences. LaBSE [16] expanded these ideas by introducing additive margin softmax, which improved the discriminative power for translation pair similarity. Despite their effectiveness, these methods often require extensive computational resources and large-scale parallel data, making them less accessible for low-resource languages.
@@
Multilingual pre-trained language models have also made significant strides in cross-lingual understanding. Cross-lingual language models (XLM) [17] introduced the objective of translation language modeling (TLM), extending masked language modeling to leverage parallel data and align representations between languages. Building on this, XLM-RoBERTa (XLM-R) [5] scaled these methods to 100 languages by pretraining on more than two terabytes of Common Crawl data [18], significantly improving performance on benchmarks for low-resource languages. XLM-R’s success underscored the importance of large-scale data and careful language sampling strategies, particularly for balancing performance across high- and low-resource languages.
@@
mT5 [6], which extended the T5 framework [19] to multilingual contexts, introduced further innovations. By adopting a text-to-text paradigm and pre-training using a large multilingual dataset covering 101 languages, mT5 achieved state-of-the-art results on various multilingual tasks. Balanced training was ensured by methods such as language sampling with controlled scaling and the use of large vocabularies. In conjunction with a span-corruption pretraining objective, the model’s encoder-decoder design enabled flexibility across tasks ranging from classification to sequence generation.
@@
Recent efforts, including EMS [20] and mSimCSE [21], have emphasized efficiency by reducing the reliance on large-scale parallel datasets and employing streamlined architectures. EMS’s integration of token-level generative and sentence-level contrastive objectives exemplifies this shift, achieving a well-aligned multilingual embedding space with minimal computational overhead. Similarly, unsupervised approaches such as those used in SimCSE [22] have been adapted for multilingual scenarios, providing competitive alternatives that further diminish the dependence on supervised data. These methods collectively push the boundaries of multilingual NLP, addressing challenges such as data scarcity and computational resource constraints.
@@
Despite these advances, it is still difficult to scale these models to include underrepresented languages and optimize their performance in low-resource settings. Research into strategies such as improved sampling distributions, multilingual adversarial training, better representation of underrepresented data, and the integration of knowledge distillation continues to enhance the accessibility and efficacy of multilingual models [11].
@@ 2.2 Adversarial Training
Adversarial training has become a cornerstone technique in machine learning, especially for improving model robustness and enhancing generalization to unseen data distributions. Initially introduced in the context of adversarial examples for image classification [8], adversarial training has since been adapted to a variety of tasks, such as named entity recognition [23], text classification [24], and multiword expression detection [25] in NLP. A central idea in adversarial training is the minimax optimization framework, which pits a discriminator against a feature extractor to encourage learning task-relevant and domain-invariant features. One influential approach involves the incorporation of gradient reversal layers to align feature spaces between domains without requiring explicit labels in the target domain. Techniques such as the domain-adversarial neural network [26] exemplify this principle, achieving significant success in cross-domain applications [27].
@@
Loss reversal [28] is a novel variation of the adversarial training paradigm that has demonstrated improved performance for domain-specific tasks. Instead of reversing the gradient at the input to the feature extractor, this method simply flips the sign of the domain discriminator’s loss before optimization. As proposed in conjunction with the HistNERo dataset for historical NER in Romanian, loss reversal encourages the model to produce more distinct embeddings between domains, boosting performance on the task. Empirical results show that integrating loss reversal with pre-trained Romanian language models, namely BERT-ro [29] and DistilRoBERT [30], leads to substantial gains in strict F1-scores, with increases of over 10% compared to the baseline non-adversarial approaches.
@@
In addition to classical adversarial setups, recent research has explored integrating adversarial training with multitask learning and self-supervised pre-training [31, 32]. These hybrid models aim to simultaneously optimize for multiple objectives, such as predicting token-level labels and aligning domain embeddings. For instance, topic-aware domain adaptation frameworks have incorporated latent topic modeling into adversarial setups, allowing models to learn domain-invariant features and topic-specific representations jointly. This combination has proven particularly effective for scientific keyphrase identification [33], where incorporating adversarial examples further enhances performance. Such approaches demonstrate the versatility of adversarial techniques in addressing domain-specific challenges across a wide range of tasks.
@@
Moreover, adversarial training has expanded beyond text classification and NER to address multilingual and cross-lingual tasks [34, 35]. In these applications, adversarial discriminators are designed to promote language-invariant representations, enabling zero-shot or few-shot learning for low-resource languages. In addition, techniques such as lateral inhibition layers have been integrated with adversarial setups to enhance the separation of embeddings for specific tasks such as multi-word expression identification [9]. These advancements highlight the ongoing evolution of adversarial training, with each iteration introducing improved mechanisms to control the interaction between feature extraction and task-specific optimization.
@@ 3.1 Training Data
The UniBERT models were pre-trained on the latest dump of the Wikipedia corpus [36], which was thoroughly filtered to ensure a high-quality dataset by employing the following heuristics:
@@
  • Removed all samples with less than five tokens using the multilingual BERT (mBERT) [2] tokenizer.
The resulting dataset contained 82GB of textual data, with the English language being the most representative (i.e., 9.8GB of data) and the Swati language the least representative (i.e., 802K of data), as depicted in Figure 1. We also measured the average number of tokens in the dataset using the mBERT tokenizer and found that each sample in the dataset had approximately 110 tokens on average, with the maximum number of tokens being approximately 10K tokens [^4] This number is truncated during pre-training and fine-tuning depending on the model maximum context window..
@@ 3.2 Model Architecture
Table 1: Comparison of our UniBERT models and the other multilingual BERT-based models available in the literature regarding their number of layers, number of hidden units on each layer, number of heads, vocabulary dimension, number of parameters, and model size.
@@
Distil-mBERT-base [37] 6 768 12 120K 135M 543M
mBERT-base [2] 12 768 12 120K 179M 715M
XLM-RoBERTa-base [17] 12 768 12 251K 279M 1.12G
XLM-RoBERTa-large [17] 24 1024 16 251K 560M 2.24G
@@
Unlike prominent multilingual models such as Distil-mBERT-base [37], mBERT-base [2], and the XLM-RoBERTa series [17], our UniBERT models offer a streamlined architecture. Table 1 shows that the UniBERT-Small model includes 6 layers with 512 hidden units per layer and 8 attention heads, which make it comparable in layer count to Distil-mBERT-base but with fewer hidden units and heads, resulting in a reduced number of parameters (87M versus 135M) and a smaller model size (334M versus 543M). This reduction is achieved while maintaining a similar vocabulary dimension of 120K, indicating that UniBERT-Small retains sufficient linguistic flexibility for multilingual tasks.
The UniBERT-XSmall and UniBERT-XXSmall configurations further emphasize parameter efficiency. UniBERT-XSmall, with 4 layers, 256 hidden units per layer, and 8 heads, holds only 39M parameters and has a model size of 148M, nearly a quarter of the size of Distil-mBERT-base. UniBERT-XXSmall, the most compact variant, is constructed with 4 layers, 128 hidden units, and 4 heads, resulting in just 19M parameters and a model size of 74M. This reduction in architectural complexity allows UniBERT-XXSmall to be highly resource-efficient, suiting applications with limited computational resources.
@@ 3.3 Training Process
Figure 2 shows the general pre-training methodology. The training framework involves a novel approach to train a smaller, more efficient BERT-based language model, called UniBERT, through three main techniques: (1) masked language modeling as for the mBERT model, (2) adversarial training using a language discriminator [38] with the loss reversal technique [28], and (3) knowledge distillation with the mBERT model serving as the teacher model.
@@ 3.3.1 Masked Language Modeling
Masked language modeling (MLM) [2] is a core component of BERT-based models that enables them to learn bidirectional representations of language. In MLM, a percentage of tokens within the input sequence are masked randomly, and the model is trained to predict these masked tokens based on the surrounding context. Formally, let $x=(x_{1},x_{2},x_{3},...,x_{n})$ be an input sequence of tokens, where $x_{i}$ represents each token. A subset of these tokens, called $x_{masked}\subset x$, is replaced by a special token [MASK], and the model optimizes the probability of recovering the original tokens $P(x_{masked}|x_{\setminus masked})$ given the context. This choice can be achieved by minimizing the cross-entropy loss over the masked tokens, as shown in Equation 1:

### v3

@@ 1 Introduction
Natural language processing (NLP) has advanced in recent years due to breakthroughs in deep learning and Transformer-based architectures (vaswani2017attention, ). Models such as Bidirectional Encoder Representations from Transformers (BERT) (devlin-etal-2019-bert, ) and its successors (liu2019roberta, ; clark2020electra, ) have redefined the state of the art in tasks such as machine translation, sentiment analysis, summarization, and question answering by learning contextualized representations that capture both syntactic and semantic nuances. This evolution has not only elevated task-specific performance, but also opened new avenues for exploring the complexities of human language.
@@
At the same time, there has been a surge of interest in developing multilingual systems capable of processing and understanding multiple languages simultaneously conneau-etal-2020-unsupervised ; xue-etal-2021-mt5 ; yano2024multilingual . Multilingual training leverages shared linguistic features across diverse languages, a benefit that is especially pronounced in low-resource settings. In addition, adversarial training 43405 has emerged as a powerful mechanism to encourage the learning of language-invariant representations, thus reducing language-specific biases and improving cross-lingual generalization in a variety of tasks avram2023multilingual ; avram2025rolargesum .
@@
However, developing robust multilingual models poses several challenges, including managing heterogeneous data distributions, accommodating language-specific nuances, and ensuring scalability without compromising performance. Adversarial training addresses these difficulties by introducing a competitive dynamic that forces the model to focus on features common to all languages. This approach not only boosts overall performance but also diminishes the model’s reliance on language-specific signals to enable better generalization across many linguistic domains xu2024survey .
@@
In this work, we introduce UniBERT, a compact and efficient BERT-based model specifically designed for multilingual processing, available in three open-source versions: (1) UniBERT-Small [^1] https://huggingface.co/avramandrei/unibert-small., (2) UniBERT-XSmall [^2] https://huggingface.co/avramandrei/unibert-xsmall., and (3) UniBERT-XXSmall [^3] https://huggingface.co/avramandrei/unibert-xxsmall.. UniBERT was pre-trained on a curated Wikipedia corpus spanning 107 languages and combines three key training components: masked language modeling devlin-etal-2019-bert , adversarial training, and knowledge distillation hinton2015distilling . This integrated methodology enables the model to achieve robust performance while significantly reducing the parameter count relative to larger models, making it an attractive solution for resource-constrained applications.
@@
We perform extensive evaluations of UniBERT on four NLP tasks: (1) named entity recognition, (2) natural language inference, (3) question answering, and (4) semantic textual similarity. The results demonstrate that our models, fine-tuned jointly on all languages, obtain better results than when they are fine-tuned on each language individually. Concretely, on average, UniBERT-Small improves the performance from 47.10% to 48.87%, UniBERT-XSmall from 40.20% to 44.35%, and UniBERT-XXSmall from 34.51% to 37.64% when trained jointly in all languages, yielding an average relative improvement of 7.72%. In contrast, larger language models show an improvement of 1.12%, underscoring the significant benefits of the multilingual joint training approach enhanced by adversarial objectives.
@@
We also performed a comprehensive statistical comparative analysis to confirm the strengths of our approach. The student’s t-test yielded a p-value of 0.0184, demonstrating that gains in UniBERT performance are significant at a significance threshold of $\alpha=0.05$. It also shows that the combination of multilingual joint training and adversarial training effectively compensates for the reduced capacity of compact models. Finally, we demonstrate the domain adaptability of UniBERT models through a comprehensive evaluation of legal documents, showing that our adversarial training approach facilitates effective knowledge transfer to specialized domains. In particular, the evaluation of the Multilingual Anonymisation for Public Administrations (MAPA) dataset DeGibertBonet2022 reveals remarkable improvements with relative gains of up to 55.71% when using multilingual joint training, further validating the effectiveness of our approach for cross-domain applications.
@@
  • We show that multilingual joint training, combined with an adversarial objective, significantly enhances cross-lingual performance, particularly in resource-constrained settings.
@@
  • We demonstrate the domain adaptability of UniBERT models through a comprehensive evaluation of legal documents, showing that our adversarial training approach facilitates effective knowledge transfer to specialized domains.
@@ 2.1 Multilingual Language Models
The development of multilingual sentence embeddings has evolved from initial methods such as LASER chaudhary2019low , which leveraged a shared long short-term memory gers2000learning encoder-decoder framework to align sentence representations across languages using parallel data. Subsequent models, such as SBERT-distill reimers2020making , utilized knowledge distillation to transfer monolingual sentence embedding capabilities into a multilingual space, allowing robust performance with fewer parallel sentences. LaBSE artetxe2019massively expanded these ideas by introducing additive margin softmax, which improved the discriminative power for translation pair similarity. Despite their effectiveness, these methods often require extensive computational resources and large-scale parallel data, making them less accessible for low-resource languages.
@@
Multilingual pre-trained language models have also made significant strides in cross-lingual understanding. Cross-lingual language models (XLM) conneau2019cross introduced the objective of translation language modeling (TLM), extending masked language modeling to leverage parallel data and align representations between languages. Building on this, XLM-RoBERTa (XLM-R) conneau-etal-2020-unsupervised scaled these methods to 100 languages by pretraining on more than two terabytes of Common Crawl data wenzek-etal-2020-ccnet , significantly improving performance on benchmarks for low-resource languages. XLM-R’s success underscored the importance of large-scale data and careful language sampling strategies, particularly for balancing performance across high- and low-resource languages.
@@
mT5 xue-etal-2021-mt5 , which extended the T5 framework raffel2020exploring to multilingual contexts, introduced further innovations. By adopting a text-to-text paradigm and pre-training using a large multilingual dataset covering 101 languages, mT5 achieved state-of-the-art results on various multilingual tasks. Balanced training was ensured by methods such as language sampling with controlled scaling and the use of large vocabularies. In conjunction with a span-corruption pretraining objective, the model’s encoder-decoder design enabled flexibility across tasks ranging from classification to sequence generation.
@@
Recent efforts, including EMS mao2024ems and mSimCSE wang2022english , have emphasized efficiency by reducing the reliance on large-scale parallel datasets and employing streamlined architectures. EMS’s integration of token-level generative and sentence-level contrastive objectives exemplifies this shift, achieving a well-aligned multilingual embedding space with minimal computational overhead. Similarly, unsupervised approaches such as those used in SimCSE gao2021simcse have been adapted for multilingual scenarios, providing competitive alternatives that further diminish the dependence on supervised data. In addition, further advances in the understanding of spoken language have emerged through multilingual mixture attention frameworks with adversarial training that adaptively capture unambiguous representations from aligned multilingual words while improving the scalability of the model for intent detection and slot filling tasks zhang2024multilingual . These methods collectively push the boundaries of multilingual NLP, addressing challenges such as data scarcity and computational resource constraints.
@@
Despite these advances, it is still difficult to scale these models to include underrepresented languages and optimize their performance in low-resource settings. Research on strategies such as improved sampling distributions, multilingual adversarial training, better representation of underrepresented data, and the integration of knowledge distillation continues to enhance the accessibility and efficacy of multilingual models xu2024survey .
@@ 2.2 Adversarial Training
Adversarial training has become a cornerstone technique in machine learning, especially for improving model robustness and enhancing generalization to unseen data distributions. Initially introduced in the context of adversarial examples for image classification 43405 , adversarial training has since been adapted to a variety of tasks, such as named entity recognition cai2024atbbc , text classification (lina2024adversarial, ; zhou2022rule, ), and multiword expression detection avram2023romanian in NLP. A central idea in adversarial training is the minimax optimization framework, which pits a discriminator against a feature extractor to encourage learning task-relevant and domain-invariant features. An influential approach involves incorporating gradient reversal layers to align feature spaces between domains without requiring explicit labels in the target domain. Techniques such as the domain-adversarial neural network ganin2016domain exemplify this principle, achieving significant success in cross-domain applications sicilia2023domain .
@@
Loss reversal avram2024histnero is a novel variation of the adversarial training paradigm that has demonstrated improved performance for domain-specific tasks. Instead of reversing the gradient at the input to the feature extractor, this method simply flips the sign of the domain discriminator’s loss before optimization. As proposed in conjunction with the HistNERo dataset for historical NER in Romanian, loss reversal encourages the model to produce more distinct embeddings between domains, boosting performance on the task. Empirical results show that integrating loss reversal with pre-trained Romanian language models, namely BERT-ro dumitrescu2020birth and DistilRoBERT avram2022distilling , leads to substantial gains in strict F1-scores, with increases of over 10% compared to the baseline non-adversarial approaches.
@@
In addition to classical adversarial setups, recent research has explored integrating adversarial training with multitask learning and self-supervised pre-training smuadu2022legal ; meftah2020multi . These hybrid models aim to simultaneously optimize for multiple objectives, such as predicting token-level labels and aligning domain embeddings. For instance, topic-aware domain adaptation frameworks have incorporated latent topic modeling into adversarial setups, allowing models to learn domain-invariant features and topic-specific representations jointly. This combination has proven particularly effective for scientific keyphrase identification smuadu2023ta , where incorporating adversarial examples further enhances performance. Such approaches demonstrate the versatility of adversarial techniques in addressing domain-specific challenges across a wide range of tasks.
@@
Moreover, adversarial training has expanded beyond text classification and NER to address multilingual and cross-lingual tasks le2024lampat ; zhang2024multilingual . In these applications, adversarial discriminators are designed to promote language-invariant representations, enabling zero-shot or few-shot learning for low-resource languages. In addition, techniques such as lateral inhibition layers have been integrated with adversarial setups to enhance the separation of embeddings for specific tasks such as multi-word expression identification avram2023multilingual . These advancements highlight the ongoing evolution of adversarial training, with each iteration introducing improved mechanisms to control the interaction between feature extraction and task-specific optimization.
@@ 3.1 Training Data
The UniBERT models were pre-trained on the latest dump of the Wikipedia corpus wikidump , which was thoroughly filtered to ensure a high-quality dataset by employing the following heuristics:
@@
  • Removed all samples with less than five tokens using the multilingual BERT (mBERT) devlin-etal-2019-bert tokenizer.
The resulting dataset contained 82GB of textual data, with the English language being the most representative (i.e., 9.8GB of data) and the Swati language the least representative (i.e., 802K of data), as depicted in Figure 1. We also measured the average number of tokens in the dataset using the mBERT tokenizer. We found that each sample in the dataset had approximately 110 tokens on average, with the maximum number of tokens being approximately 10K tokens [^4] This number is truncated during pre-training and fine-tuning depending on the model’s maximum context window..
@@ 3.2 Model Architecture
Table 1: Comparison of our UniBERT models and the other multilingual BERT-based models available in the literature regarding their number of layers, number of hidden units in each layer, number of heads, vocabulary dimension, number of parameters, and model size.
@@
Distil-mBERT-base Sanh2019DistilBERTAD 6 768 12 120K 135M 543M
mBERT-base devlin-etal-2019-bert 12 768 12 120K 179M 715M
XLM-RoBERTa-base conneau2019cross 12 768 12 251K 279M 1.12G
XLM-RoBERTa-large conneau2019cross 24 1024 16 251K 560M 2.24G
XLM-V-base liang2023xlm 12 768 12 901K 778M 3.12G
@@
Unlike prominent multilingual models such as Distil-mBERT-base Sanh2019DistilBERTAD , mBERT-base devlin-etal-2019-bert , and the XLM-RoBERTa series conneau2019cross , our UniBERT models offer a streamlined architecture. Table 1 shows that the UniBERT-Small model includes 6 layers with 512 hidden units per layer and 8 attention heads, which make it comparable in layer count to Distil-mBERT-base but with fewer hidden units and heads, resulting in a reduced number of parameters (87M versus 135M) and a smaller model size (334M versus 543M). This reduction is achieved while maintaining a similar vocabulary dimension of 120K, indicating that UniBERT-Small retains sufficient linguistic flexibility for multilingual tasks.
Further comparison of our UniBERT models with the XLM family reveals significant efficiency differences. While XLM-RoBERTa-base (i.e., 279M parameters, 1.12G size), XLM-RoBERTa-large (i.e., 560M parameters, 2.24G size), and XLM-V-base (i.e., 778M parameters, 3.12G size) liang2023xlm rely on increasingly larger architectures with extensive vocabularies, UniBERT-Small reduces the model size to less than 12% of XLM-V-base’s footprint. This efficiency gap widens even more with UniBERT-XSmall (39M parameters) and UniBERT-XXSmall (19M parameters), the latter being approximately 30 times smaller than XLM-RoBERTa-large and 40 times smaller than XLM-V-base, while maintaining effective multilingual capabilities. These dramatic size reductions demonstrate our approach’s effectiveness in creating compact yet functional models for resource-constrained environments.
@@ 3.3 Training Process
Figure 2 shows the general pre-training methodology. The training framework involves a novel approach to train a smaller, more efficient BERT-based language model, called UniBERT, through three main techniques: (1) masked language modeling as for the mBERT model, (2) adversarial training using a language discriminator chen2018adversarial with the loss reversal technique avram2024histnero , and (3) knowledge distillation with the mBERT model serving as the teacher model.
@@ 3.3.1 Masked Language Modeling
Masked language modeling (MLM) devlin-etal-2019-bert is a core component of BERT-based models that enables them to learn bidirectional representations of language. In MLM, a percentage of tokens within the input sequence are masked randomly, and the model is trained to predict these masked tokens based on the surrounding context. Formally, let $x=(x_{1},x_{2},x_{3},...,x_{n})$ be an input sequence of tokens, where $x_{i}$ represents each token. A subset of these tokens, called $x_{masked}\subset x$, is replaced by a special token [MASK], and the model optimizes the probability of recovering the original tokens $P(x_{masked}|x_{\setminus masked})$ given the context. This choice can be achieved by minimizing the cross-entropy loss over the masked tokens, as shown in Equation 1: